# How Well Do LLMs Handle Cantonese? Benchmarking Cantonese Capabilities of Large Language Models

### v2

@@ Abstract
The rapid evolution of large language models (LLMs) has transformed the competitive landscape in natural language processing (NLP), particularly for English and other data-rich languages. However, underrepresented languages like Cantonese, spoken by over 85 million people, face significant development gaps, which is particularly concerning given the economic significance of the Guangdong-Hong Kong-Macau Greater Bay Area, and in substantial Cantonese-speaking populations in places like Singapore and North America. Despite its wide use, Cantonese has scant representation in NLP research, especially compared to other languages from similarly developed regions. To bridge these gaps, we outline current Cantonese NLP methods and introduce new benchmarks designed to evaluate LLM performance in factual generation, mathematical logic, complex reasoning, and general knowledge in Cantonese, which aim to advance open-source Cantonese LLM technology. We also propose future research directions and recommended models to enhance Cantonese LLM development [^1] The code and data are available anonymously github. The anonymously github link is as follow: https://github.com/jiangjyjy/Yue-Benchmark.
@@ 1 Introduction
Increasingly impactful and LLMs have emerged (e.g., GPT-X, Llama-X, etc.), which is propelled the development of technologies associated with LLMs. As shown in Figure 1, NLP research has predominantly concentrated on creating models for English and a few other languages that have substantial data resources Aji et al. (2022). The scarcity of data is often identified as the primary obstacle impeding advancements in NLP for languages that are less represented Hu et al. (2020); Joshi et al. (2020); Aji et al. (2022), particularly for LLM-related technologies.
@@ 2.1 Cantonese small-scale neural network
Sentiment Analysis. Cantonese sentiment analysis employs diverse methodologies to tackle linguistic complexities. Early approaches used Naive Bayes and SVMs with character-based bi-grams, while later studies utilized Hidden Markov Models for text segmentation and part-of-speech tagging, developing emotion-specific dictionaries via rule-based systems Zhang et al. (2011); Chen et al. (2013, 2015). More recent studies have enhanced classification accuracy using both supervised and unsupervised methods across various domains, with Lee et al. (2019) exploring fine-grained emotion analysis across languages Ngai et al. (2018); Xiang et al. (2019); Lee (2019).
@@ 3.1 Existing Cantonese data
interactions, with Cantonese becoming the main focus of most bilingual dictionaries by the 19th century Xiang et al. (2024). Hong Kong led the development of Cantonese linguistic resources, including bilingual corpora from the Legislative Council Wu (1994), a one-million-character Cantonese corpus from children’s dialogues Hun-tak Lee (1999), and specialized corpora for Cantonese-speaking children Yip and Matthews (2007). Significant contributions also came from television and theater productions Leung and Law (2001), and the University of Hong Kong’s work on spontaneous speech, focusing on transcription and tagging Ping-Wai (2006). A parallel Cantonese-Standard Chinese corpus was developed for machine translation, sourced from television broadcasts Lee (2011). Recent efforts have focused on closing the data gap between Cantonese and other major languages through a small dependency treebank and a comprehensive bilingual dictionary, enhancing tools for translation Xiang et al. (2024).
@@ 3.2 New benchmarks construction
Figure 3: Translation example.
@@ 4.4 Results and analysis
Enhance training data volume by utilizing models that excel in translating Cantonese data.
@@ 5.2 Opportunities
Large language model. Based on the analysis above, we compile Figure 5, which presents the best LLMs in 0-shot and 5-shot, and suggests LLM series for various tasks. For work related to LLMs, we recommend using newer versions of the Qwen, Mixtral, Llama, and Yi series (Table B,910B). For tasks that involve only prompting, without the need for LLM training, we also recommend using closed-source models such as GPT, GLM, and Sensechat series models.

### v3

@@ Abstract
The rapid evolution of large language models (LLMs) has transformed the competitive landscape in natural language processing (NLP), particularly for English and other data-rich languages. However, underrepresented languages like Cantonese, spoken by over 85 million people, face significant development gaps, which is particularly concerning given the economic significance of the Guangdong-Hong Kong-Macau Greater Bay Area, and in substantial Cantonese-speaking populations in places like Singapore and North America. Despite its wide use, Cantonese has scant representation in NLP research, especially compared to other languages from similarly developed regions. To bridge these gaps, we outline current Cantonese NLP methods and introduce new benchmarks designed to evaluate LLM performance in factual generation, mathematical logic, complex reasoning, and general knowledge in Cantonese, which aim to advance open-source Cantonese LLM technology. We also propose future research directions and recommended models to enhance Cantonese LLM development [^1] The code and data are available on github: https://github.com/jiangjyjy/Yue-Benchmark.
@@ 1 Introduction
Increasingly impactful and LLMs have emerged (e.g., GPT-X, Llama-X, DeepSeek-X, etc.), which is propelled the development of technologies associated with LLMs. As shown in Figure 1, NLP research has predominantly concentrated on creating models for English and a few other languages that have substantial data resources Aji et al. (2022). The scarcity of data is often identified as the primary obstacle impeding advancements in NLP for languages that are less represented Hu et al. (2020); Joshi et al. (2020); Aji et al. (2022), particularly for LLM-related technologies.
@@ 2.1 Cantonese small-scale neural network
Sentiment Analysis. Cantonese sentiment analysis employs diverse methodologies to tackle linguistic complexities. Early approaches used Naive Bayes and SVMs with character-based bi-grams, while later studies utilized Hidden Markov Models for text segmentation and part-of-speech tagging, developing emotion-specific dictionaries via rule-based systems Zhang et al. (2011); Chen et al. (2013, 2015). More recent studies have enhanced classification accuracy using both supervised and unsupervised methods across various domains, with Lee (2019) exploring fine-grained emotion analysis across languages Ngai et al. (2018); Xiang et al. (2019); Lee (2019).
@@ 3.1 Existing Cantonese data
The documentation of dialects expanded due to trade and cultural interactions, with Cantonese becoming the main focus of most bilingual dictionaries by the 19th century Xiang et al. (2024). Hong Kong led the development of Cantonese linguistic resources, including bilingual corpora from the Legislative Council Wu (1994), a one-million-character Cantonese corpus from children’s dialogues Hun-tak Lee (1999), and specialized corpora for Cantonese-speaking children Yip and Matthews (2007). Significant contributions also came from television and theater productions Leung and Law (2001), and the University of Hong Kong’s work on spontaneous speech, focusing on transcription and tagging Ping-Wai (2006). A parallel Cantonese-Standard Chinese corpus was developed for machine translation, sourced from television broadcasts Lee (2011). Recent efforts have focused on closing the data gap between Cantonese and other major languages through a small dependency treebank and a comprehensive bilingual dictionary, enhancing tools for translation Xiang et al. (2024).
@@ 3.2 New benchmarks construction
Figure 3: Examples in Yue-Benchmark.
@@ 4.4 Results and analysis
Enhancing Data Quality and Cost-Effectiveness for Cantonese LLMs.
@@ 5.2 Opportunities
Large language models. Based on the analysis above, we compile Figure 5, which presents the best LLMs in 0-shot and 5-shot, and suggests LLM series for various tasks. For work related to LLMs, we recommend using newer versions of the Qwen, Mixtral, Llama, and Yi series (Table B,910B). For tasks that involve only prompting, without the need for LLM training, we also recommend using closed-source models such as GPT, GLM, and Sensechat series models.
6 Conclusion and Outlook
Cantonese, spoken by over 85 million people, lags in natural language processing development, especially in large language models. To address this gap, we summarize existing Cantonese NLP methods and introduce four new benchmarks (Yue-Truthful, Yue-GSM8K, Yue-ARC-C, Yue-MMLU) and a translation dataset (Yue-TRANS). We evaluate 35 mainstream LLMs on these benchmarks, identifying current strengths and weaknesses. This work lays a foundation for advancing Cantonese LLM related technology.
Future efforts focus on building larger, high-quality Cantonese corpora and optimizing models for Cantonese-specific tasks. Collaboration among global researchers accelerates progress, helping Cantonese NLP catch up with other languages, enriching the experiences of Cantonese speakers.
@@ 6 Conclusion and Outlook
Acknowledgements
We want to thank our anonymous AC and reviewers for their feedback. This work was supported by Hong Kong Innovation and Technology Commission’s Innovation and Technology Fund (Award No. ITS/269/22FP).
@@ A.4 Cantonese large language model
Regarding the training and reasoning technologies for LLMs associated with mainstream languages, there is no development specific to Cantonese such as LoRA Hu et al. (2021); Wang et al. (2024b, a), reasoning Gao et al. (2024); Havrilla et al. (2024), etc.